OpenStudy (anonymous):

What is the order of the augmented matrix for this { x+y=3 2x−y=3 linear system?

OpenStudy (anonymous):

\[\left[\begin{matrix}x+y=3 &\\ 2x-y=3& \end{matrix}\right]\]

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

2 × 2 3 × 3 2 × 3 3 × 2

OpenStudy (owlcoffee):

The Augmented matrix of any system of equalities is always represented by the coeficents of the variables and the constante they represent, say for instance I have the system of equation: \[Ax+By+Cz=K_1\] \[A'x+B'y+C'z=K'\] \[A''x+B''y+C''z=K''\] The "augemented matrix" will be a 4 by 3 matrix, because it'll look like this: \[\left[\begin{matrix}A & B & C &K \\ A' & B' & C' & K'\\ A'' & B'' & C'' & K''\end{matrix}\right]\]

OpenStudy (owlcoffee):

A MAtrix is always represented by first writing the number of rows and then the number of columns: "nxm"

OpenStudy (anonymous):

so its a 2by 2

OpenStudy (owlcoffee):

That's incorrect, try again.

OpenStudy (anonymous):

so its a 3by 3

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (owlcoffee):

Close, but you have it wrong on the columns

OpenStudy (anonymous):

got to be a 2 by 3

OpenStudy (owlcoffee):

You are correct.